# SAQ-SAM: Semantically-Aligned Quantization for Segment Anything Model

### v1

@@ Abstract
Segment Anything Model (SAM) exhibits remarkable zero-shot segmentation capability; however, its prohibitive computational costs make edge deployment challenging. Although post-training quantization (PTQ) offers a promising compression solution, existing methods yield unsatisfactory results when applied to SAM, owing to its specialized model components and promptable workflow: (i) The mask decoder’s attention exhibits extreme outliers, and we find that aggressive clipping (ranging down to even 100$\times$), instead of smoothing or isolation, is effective in suppressing outliers while maintaining semantic capabilities. Unfortunately, traditional metrics (e.g., MSE) fail to provide such large-scale clipping. (ii) Existing reconstruction methods potentially neglect prompts’ intention, resulting in distorted visual encodings during prompt interactions. To address the above issues, we propose SAQ-SAM in this paper, which boosts PTQ of SAM with semantic alignment. Specifically, we propose Perceptual-Consistency Clipping, which exploits attention focus overlap as clipping metric, to significantly suppress outliers. Furthermore, we propose Prompt-Aware Reconstruction, which incorporates visual-prompt interactions by leveraging cross-attention responses in mask decoder, thus facilitating alignment in both distribution and semantics. To ensure the interaction efficiency, we also introduce a layer-skipping strategy for visual tokens. Extensive experiments are conducted on different segmentation tasks and SAMs of various sizes, and the results show that the proposed SAQ-SAM consistently outperforms baselines. For example, when quantizing SAM-B to 4-bit, our method achieves 11.7% higher mAP than the baseline in instance segmentation task.
@@ 1 Introduction
Segment Anything Model [10] (SAM) shows promising applications as the base model for promptable segmentation. Benefiting from sufficient pre-training and prompt-guided fine-tuning, SAM exhibits strong zero- or few-shot generalization capability to interactively segment regions of interest to the user. However, SAM’s powerful representation capability is accompanied by a large number of parameters and high computational costs, limiting its potential on resource-constrained devices [38, 2].
@@
Figure 1: Visualization of extreme activation distributions and different clipping results. QK activations in the mask decoder show highly skewed distributions, with most data concentrated in a narrow range while outliers can exceed 180 times the normal range. MSE provides an overly wide clipping range, whereas our Perception-Consistency Clipping method can identify outliers more precisely. The activation tensors are hooked from the Query input in the mask decoder.
Figure 2: Overview of SAQ-SAM. The proposed Perceptual-Consistency Clipping method guides QK activation quantization clipping by minimizing the Attention Focus deviation from FP, thereby semantically preserving the perceptual alignment. Our Prompt-Aware Reconstruction method incorporates prompt-image interactions within the mask decoder into local reconstruction. Through minimizing the interaction response error supervised by the FP model, quantization model learns correspondence between visual features and prompt intentions, thus facilitating dual alignment at both the distributional and semantic levels.
@@
Model quantization can reduce computational overhead and model size by converting weights and activations to low-bit integers [7, 5, 4], and post-training quantization (PTQ) [26, 11], which efficiently calibrates quantization parameters using a small set of unlabeled data, stands out as a promising approach. Unfortunately, SAM features the unique activation distribution and network architecture, which pose new challenges to PTQ, rendering conventional methods inadequate. To this end, several studies have attempted to propose specific solutions for SAM’s characteristics. For instance, PQ-SAM [21] hierarchically clusters similar channels to learn unified transformation factors. However, it is limited to the image encoder with a classical Transformer structure and lacks adaptability to the mask decoder, which features more complex connection paths. In practice, although SAM’s parameters are primarily concentrated in the image encoder, the mask decoder can also introduce significant computational overhead. For example, in panoramic semantic segmentation task, given a single image, the mask decoder must perform multiple large-batch inferences for numerous prompts, whereas the image encoder only requires a single forward pass [1]. Thus, PTQ4SAM [21] proposes an improvement for quantization bottleneck in the mask decoder, which merges bimodal distributions in post-key-linear activations through QK equivalence transformation. However, the performance of these methods, which aim to better align the full-precision model (FP) at distribution level, is still far from satisfactory, especially for low-bit quantization.
In this paper, we empirically observe that the QK activation distribution of SAM’s mask decoder exhibits extreme outliers, forming a bottleneck for quantization performance. To deal with outliers, traditional methods typically employ smoothing or isolation schemes. Nevertheless, the mask decoder struggles to adapt to these methods due to its complex connection paths. Fortunately, we find an interesting phenomenon: directly clipping these outliers with an aggressive magnitude has very little impact on the final segmentation performance. For example, with SAM-B’s query input range spanning [-178,182], aggressively truncating it to [-5, 5] does not hurt any accuracy. This is because, although modifying the original distribution, it maintains the semantical alignment, as shown in Figure 3. It is well-known that quantization typically benefits from a more compact clipping range for higher resolution, for example, using the clipping range of [-5, 5] yields an 11.8% mAP improvement (from 17.4% to 29.2%) for 6-bit SAM-B. However, traditional distance metrics (e.g., MSE), which aim to align distributions, fail to provide such large-scale clipping. For example, as illustrated in Figure 1, MSE gives a clipping range of [-167,177], which results in severe degradation of accuracy (from 37.0% in FP to 17.4%). Thus, instead of the distributional level, we aim to establish a semantic-level metric that facilitates effective outlier clipping by ensuring semantic alignment. Moreover, previous quantization reconstruction methods primarily focus on local reconstruction of FP responses while neglecting the semantic intent of prompts. This oversight leads to overfitting on calibration images, thus causing feature distortions in visual-prompt interactions and ultimately compromising segmentation accuracy. To this end, we are also motivated to explore the integration of prompt interactions into the reconstruction process to facilitate semantic alignment.
@@
With the above insights, we propose SAQ-SAM, which significantly improves PTQ performance for SAM by promoting semantic-level alignment with the FP model. Specifically, we propose Perceptual-Consistency Clipping to drastically suppress outliers in QK activations, without smoothing or isolation operations. By proactively measuring the quantization-induced deviation of attention focus, the clipping range is determined without being affected by the magnitude of outliers. As illustrated in Figure 3, our method evidently mitigates the attention functional degradation issue that is obvious in distribution-level alignment methods (e.g., MSE) ,thereby maintaining semantic consistency. Furthermore, we propose Prompt-Aware Reconstruction, which learns quantization parameters under the guidance of prompts. By leveraging the mask decoder to facilitate interaction between prompts and image tokens, we reconstruct the interaction responses supervised by the FP model. This preserves the correspondence between visual features and prompts, as well as alleviating the overfitting issue of local response reconstruction methods. The main contributions are summarized as follows:
@@
  • We observe that aggressively clipping outliers in SAM has minimal impact on segmentation performance. With this insight, we propose Perceptual-Consistency Clipping, which utilizes attention focus deviation as a metric to guide large-magnitude clipping, and thus suppressing outliers while maintaining semantic alignment.
  • We propose Prompt-Aware Reconstruction, which integrates prompt-image interactions within the mask decoder into local reconstruction. This enables the quantized model to align with the FP model in both visual features and prompt intentions, thus facilitating dual alignment at the distributional and semantic levels.
@@
  • We conduct extensive experiments on SAMs of various sizes across both prompt segmentation and panoptic semantic segmentation tasks, demonstrating the consistent superiority of SAQ-SAM over baseline methods. For instance, in instance segmentation, our 4-bit SAQ-SAM achieves an average 14% mAP improvement on SAM-B and maintains nearly lossless accuracy on SAM-L.
@@ 2.1 Segment Anything
Segment Anything [10, 27] (SAM) has emerged as a versatile and promptable image segmentation model. Trained on an extensive dataset comprising over 11 million images and 1 billion masks, SAM demonstrates remarkable generalization across various downstream applications, including edge detection, object proposal, and instance segmentation. Its capabilities extend to critical domains such as medical image annotation [24, 3], where it shows significant promise as a powerful diagnostic support tool. However, SAM’s powerful capabilities come with extensive memory and computational costs. Although some studies have proposed various lightweight variants to accelerate inference [39, 36, 35], the model parameters of these methods are still kept in floating-point type. This prevents them from benefiting from efficient low-precision integer operation units, resulting in suboptimal efficiency.
2.2 Post-Training Quantization
@@
Model quantization [7, 15, 14] reduces memory and computation costs by converting floating-point values to integers. To mitigate performance degradation, various works were proposed, which are broadly categorized into two types: Quantization-Aware Training (QAT) and PTQ. QAT fine-tunes quantized models with labeled data at high cost [8, 13], while PTQ efficiently uses a handful of unlabeled samples to set quantization parameters [25, 18].
@@
SAM utilizes Transformer structures for both its image encoder and mask decoder. Various works have dedicated to developing PTQ methods for Transformer structures. SmoothQuant [33] employs equivalent transformations to smooth the activation distribution, mitigating the impact of harmful outliers. OmniQuant [30] improves the applicability of smoothing methods by making equivalent transformation factors and clipping ranges learnable. RepQ-ViT [16] and RepQuant [17] decouple quantization from inference, bridging the gap through scale reparameterization for efficient hardware-friendly quantization. Besides these distribution-adjustment methods, some methods exploit the characteristics of the attention mechanism. For example, PTQ-ViT [22] determines quantization clipping by maintaining attention rank consistency and search mixed-precision scheme using the nuclear norm of module response. Other methods, such as BRECQ [11], QDrop [32], and PD-Quant [20], aim to learn optimal quantization parameters. They reconstruct the internal response of the quantized model under the supervision of the FP model.
@@
Despite the aforementioned methods have achieved remarkable performance in dealing with Transformer structures, SAM’s distinctive activation distribution poses new challenges, leaving their performance subpar. To this end, several works try to make improvements. To tackle the channel-wise distribution imbalance in activations, PQ-SAM [21] hierarchically clusters channels with similar distributions and learns shared transformation factors for each group, thereby reducing optimization complexity. PTQ4SAM [23] identify and integrate the bimodal distribution in post-key-linear activation of SAM’s mask decoder, while implementing an adaptive granularity quantizer for post-softmax activation to adapt to different attention module. However, these distribution-level improvements have shown unsatisfactory performance in practice.
@@ 3.1 Preliminaries
Transformer Layers in SAM: In SAM’s image encoder, the Transformer layers employ window attention and global attention alternately [12], and the former divides the image into non-overlapping windows and computes self-attention within each window without shifting. In the mask decoder, a lightweight Two-Way Transformer is utilized to update both the image embedding and prompt tokens via cross-attention, facilitating the information interaction between prompt and image. Specifically, the token-to-image cross-attention utilizes prompt tokens as queries and employs image tokens as keys and values, whereas the image-to-token cross-attention implements the inverse configuration to ensure comprehensive feature interaction.
@@
[…]
@@
The process of determining clipping boundaries is crucial for improving the quantization performance in PTQ, which is also called calibration.
@@ 3.2 Perceptual-Consistency Clipping
Insight. With the observation that aggressively clipping the extreme outlier in QK activations, although significantly modifying the distribution, doesn’t hurt the segmentation performance, we aim to break through the limitations of distribution alignment and turn to leverage the semantic nature of attention mechanisms for potential breakthroughs. As state in [31], in Transformers architecture, the attention mechanism captures the semantic information perceived by the model. It allocates greater focus to regions of interest, thereby reinforcing the features most relevant to the task. Building on this, we seek to leverage the potential relationships modeled by attention to preserve semantic-level consistency before and after quantization.
Figure 3: Comparison of attention heatmaps using different outlier clipping methods. The distribution-aligned MSE method leads to significant attention degradation, whereas our sematic-aligned PCC method maintains consistency with the FP model. The heatmaps are generated from the prompt-query cross-attention module in the first layer of the mask decoder.
Attention Focus Overlap metric. Inspired by this, we define high-attention tokens as perceptual focus and maximize the overlap of the focus regions in the attention maps before and after quantization, thereby maintaining consistency in attention perception. The procedure of the metric formulation is described as follows.
@@
Given the above attention scores, we aim to distinguish salient regions and define them as the Attention Focus. To achieve this, threshold factor $\theta\in(0,1)$ is used to filter the significant values, resulting in a binarized Attention Focus Mask as follows:
$M_{A}=M(A_{w})=\mathbbm{1}\left\{A_{w}>\theta\cdot\max\left(A_{w}\right)\right\}\in\mathbb{R}^{N_{q}\times N_{k}},$ (6)
where $\mathbbm{1}\left\{\cdot\right\}$ represents an indicator function that returns 1 if the condition is true, otherwise 0.
@@
where the variables with a hat superscript ($\hat{A}_{w}$/$\hat{M}_{A}$) denote the scores/masks from the quantized model. This metric lies in the range (0,1), and the distance function for Perceptual-Consistency Clipping is defined as:
@@
Then, we leverage the defined Attention Focus Overlap metric to determine the optimal clipping boundaries (i.e. $x_{low}$ and $x_{up}$) for QK activations. Unlike conventional methods constrained by distributional limitations, our proposed clipping metric operates at a semantic level, enabling more effective clipping decisions. As shown in Figure 3, while MSE-based calibration causes severe attention degradation, our clipping method achieves closer results to the FP model, efficiently preserves the perceptual capabilities of the quantized attention module.
Note that the proposed metric is specifically applied to QK activations, while other activations continue to be calibrated using the MSE metric. This strategy enables a balanced trade-off between distributional alignment and semantic fidelity across the model. By maintaining this balance, we prevent the risk of overemphasizing semantic alignment, which could otherwise lead to distributional collapse and degrade overall model performance.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Segment Anything Model (kirillov2023segment) (SAM) shows promising applications as the base model for promptable segmentation. Benefiting from sufficient pre-training and prompt-guided fine-tuning, SAM exhibits strong zero- or few-shot generalization capability to interactively segment regions of interest to the user. However, SAM’s powerful representation capability is accompanied by a large number of parameters and high computational costs, limiting its potential on resource-constrained devices (zhang2024efficientvit; chen2023slimsam).
@@
Figure 1: Visualization of extreme activation distributions in mask decoder and the performance of different clipping methods. QK activations in the mask decoder show highly skewed distributions, with most data concentrated in a narrow range while outliers can exceed 180 times the normal range. MSE provides an overly wide clipping range, whereas our Perceptual-Consistency Clipping (PCC) method can identify outliers more precisely.
@@
Model quantization (li2024htq; liu2024eda; xiao2024binaryvit) can reduce computational overhead and model size by converting weights and activations to low-bit integers (li2023vit; gholami2021survey; choi2018pact), and post-training quantization (PTQ) (adaround; li2023psaq), which efficiently calibrates quantization parameters using a small set of unlabeled data, stands out as a promising approach. Unfortunately, SAM features the unique activation distribution and network architecture, which pose new challenges to PTQ, rendering conventional methods inadequate. To this end, several studies (li2024privacy) have attempted to propose specific solutions for SAM’s characteristics. For instance, PQ-SAM (pq-sam) hierarchically clusters similar channels to learn unified transformation factors. However, it is limited to the image encoder with a classical Transformer structure and lacks adaptability to the mask decoder, which features more complex connection paths. In practice, although SAM’s parameters are primarily concentrated in the image encoder, the mask decoder can also introduce significant computational overhead. As shown in Figure 2, in semantic segmentation task, given a single image, the mask decoder must perform multiple large-batch inferences for numerous prompting points, whereas the image encoder requires much fewer forward passes (chen2023semantic). To this end, PTQ4SAM (ptq4sam) is proposed for quantization bottleneck in the mask decoder, which merges bimodal distributions in post-key-linear activations through QK weight transformation. However, it still seeks to align full-precision (FP) model at distribution level, leading to unsatisfactory performance, especially for low-bit quantization.
Figure 2: Comparison of total inference time between image encoder and mask decoder in semantic segmentation task.
In this paper, we comprehensively investigate SAM’s model architecture and interactivity in promptable segmentation, and find that the semantic misalignment constitutes the main challenge for low-bit quantization. In the following, we will discuss the misalignment issues in two critical techniques of PTQ: outlier handling and reconstruction.
i) Semantic Misalignment in Outlier Handling. QK activations in SAM’s mask decoder exhibits extreme outliers, as shown in Figure 1(a) and Appendix B.2. Fortunately, we find an interesting phenomenon: aggressively clipping these outliers has little impact on final segmentation performance. As shown in Figure 1(b), clipping the skewed distributed activation from [-167, 177] and [-114, 54] to [-1, 1] has little impact. More importantly, such large-scale range reduction improves quantization resolution, thereby reducing quantization errors. To this end, we perform a grid search of large-scale clipping intervals for quantization calibration, as shown in Figure 1(c), where [-1, 1] achieves 12.4% performance improvement compared with traditional MSE metric. Thus, distribution-based metrics are no longer applicable to segmentation tasks, resulting in severe semantic degradation, as shown in Figure 3. As a result, it is crucial to establish a semantic level metric to drive optimal-scale clipping.
Figure 3: Attention heatmaps in the mask decoder with different quantization clipping methods. The distribution-aligned MSE leads to significant attention degradation, whereas our semantic-aligned PCC maintains the consistency with the FP model.
ii) Semantic Misalignment in Reconstruction. Quantization reconstruction is a prevalent method for restoring model performance. However, existing methods typically target at simple tasks with single branch model architecture, thus exhibiting suboptimal performance in promptable segmentation task with encoder-decoder architecture. This is because, locally reconstructing the FP responses in the image encoder overlooks the semantic intent of prompts, thus causing feature distortions in visual-prompt interactions and ultimately compromising segmentation accuracy. To this end, to ensure quantization accuracy, it is crucial to facilitate semantic alignment during reconstruction.
@@
With the above insights, we propose SAQ-SAM, which significantly improves PTQ performance for SAM by promoting semantic-level alignment with the FP model. Specifically, for outlier suppression, we propose Perceptual-Consistency Clipping (PCC) for QK activations. By proactively measuring the quantization-induced deviation in attention focus, PCC achieves magnitude-independent outlier suppression with semantic alignment. As illustrated in Figure 3, PCC evidently mitigates the attention functional degradation issue. Further, we propose Prompt-Aware Reconstruction (PAR), which learns quantization parameters under the guidance of prompts. By leveraging SAM’s off-the-shell mask decoder to facilitate interaction between prompts and image tokens, PAR reconstructs the interaction responses supervised by the FP model. This preserves the correspondence between visual features and prompts, thus facilitating alignment in both distribution and semantic. The main contributions are summarized as follows:
@@
  • We observe that aggressively clipping outliers in SAM has minimal impact on segmentation performance. With this insight, we propose Perceptual-Consistency Clipping, which utilizes attention focus deviation as a metric to guide large-scale clipping, and thus suppressing outliers while maintaining semantic alignment.
  • We propose Prompt-Aware Reconstruction, which integrates image-prompt interactions into reconstruction. This enables the quantization model to align with the FP model in both feature modeling and prompt following. We also design a layer-skipping strategy of image tokens to ensure interaction efficiency.
@@
  • We conduct extensive experiments on SAMs of various sizes across three mainstream tasks, demonstrating the consistent superiority over baseline methods. For instance, in instance segmentation, our 4-bit SAQ-SAM achieves an average 14% mAP improvement on SAM-B and maintains nearly lossless accuracy on SAM-L.
Figure 4: Overview of SAQ-SAM. The proposed PCC guides quantization clipping of QK activation by minimizing the Attention Focus deviation from FP, thereby semantically preserving the perceptual alignment. Our PAR incorporates image-prompt interactions into per-stage reconstruction, utilizing the off-the-shell module in the mask decoder. Through minimizing the interaction response error supervised by the FP model, quantization model learns correspondence between visual features and prompt intentions, thus facilitating dual alignment at both the distributional and semantic levels.
@@ 2 Related Work
Segment Anything. Segment Anything (SAM) (kirillov2023segment; ravi2024sam) has emerged as a versatile and promptable image segmentation tool. Pre-trained on an extensive dataset, SAM demonstrates remarkable generalization across various downstream tasks, including instance segmentation (chen2024rsprompter), semantic segmentation (chen2023semantic), and oriented object detection (yu2023h2rbox). Its capabilities extend to critical domains such as medical image annotation (MedSAM; sammed2d), where it shows significant promise as a powerful diagnostic support tool. Thanks to its excellent generalization capabilities, SAM is also be used as an auxiliary branch for industrial anomaly detection (li2025sam_ano). However, SAM’s powerful capabilities come with extensive memory and computational costs. Although some studies have proposed various lightweight variants to accelerate inference (fastsam; mobilesam; efficientsam), the model parameters of these methods are still kept in floating-point type. This prevents them from benefiting from efficient low-precision integer operation units, resulting in suboptimal efficiency.
@@
Post-Training Quantization. Model quantization (li2023qft; li2022patch; liu2025cachequant) reduces memory and computation costs by converting floating-point values to integers. To mitigate the performance degradation caused by precision loss, researchers have made various efforts, among which PTQ efficiently uses a handful of unlabeled samples to set quantization parameters (lin2024awq; liu2024dilatequant; li2022dual).
@@
The backbone of SAM is built on the Transformer architecture, a structure for which PTQ have been extensively researched and developed. SmoothQuant (smoothquant) employs equivalent transformations to smooth the activation distribution, mitigating the impact of harmful outliers. OmniQuant (omniquant) improves the applicability of smoothing methods by learning equivalent transformation factors and clipping ranges. RepQ-ViT (repq-vit) and RepQuant (repquant) decouple quantization from inference, bridging the gap through scale reparameterization for efficient hardware-friendly quantization. In addition to these distribution-adjustment methods, some methods exploit the characteristics of the attention mechanism. For example, PTQ-ViT (ptqvit) determines quantization clipping by maintaining the consistency of the attention rank. Other methods, such as BRECQ (brecq), QDrop (qdrop), and PD-Quant (pd), aim to learn optimal quantization parameters. They reconstruct the internal response of the quantization model under the supervision of the FP model.
@@
Despite the aforementioned methods achieved remarkable performance in dealing with classical Transformer structures, SAM’s distinctive activation distribution poses new challenges, leaving their performance subpar. To this end, several works try to make improvements. To tackle the channel-wise distribution imbalance in activations, PQ-SAM (pq-sam) hierarchically clusters channels with similar distributions and learns shared transformation factors for each group, thereby reducing optimization complexity. PTQ4SAM (ptq4sam) identify and integrate the bimodal distribution in post-key-linear activation in mask decoder, and implement an adaptive granularity quantizer for post-softmax activation to adapt to different types of attention module. However, these distribution-level improvements have shown unsatisfactory performance in practice.
@@ 3.1 Preliminaries
Transformer Layers in SAM: In SAM’s image encoder, the Transformer layers employ window attention and global attention alternately (win_attn), and the former divides the image into non-overlapping windows and computes self-attention within each window without shifting. In the mask decoder, a lightweight Two-Way Transformer is utilized to update both the image embedding and prompt tokens via cross-attention, facilitating the information interaction between prompt and image. Specifically, the token-to-image cross-attention utilizes prompt tokens as queries and employs image tokens as keys and values, whereas the image-to-token cross-attention implements the inverse configuration to ensure comprehensive feature interaction.
@@
where $x$ and $x_{q}$ are the floating-point and quantized values, respectively. The dequantized value $\hat{x}$ approximates $x$. $\left\lfloor\cdot\right\rceil$ is the round-to-nearest operation. clip function truncates values outside the $b$-bit range. The quantization scale $s$ and zero-point $z$ are PTQ parameters to be searched, determined by the clipping boundaries $x_{low}$ and $x_{up}$ as follows:
@@
The process of determining clipping boundaries is crucial for PTQ performance, which is also called calibration.
@@ 3.2 Perceptual-Consistency Clipping
Insight. With the observation that aggressively clipping the extreme outlier in QK activations, although significantly modifying the distribution, doesn’t hurt the segmentation performance, we aim to break through the limitations of distribution alignment and turn to leverage the semantic nature of attention mechanisms for PTQ calibration. As stated in (vaswani2017attention), in Transformers architecture, the attention mechanism captures the semantic information perceived by the model. It allocates greater focus to regions of interest, thereby reinforcing the features most relevant to the task. Building on this, it’s reasonable to leverage the potential relationships modeled by attention to preserve semantic-level consistency before and after quantization.
Attention Focus Overlap Metric. Inspired by above insights, we define highly attented region as the perceptual focus and maximize the overlap of the perceptual focus before and after quantization, thereby maintaining consistency in attention perception. The procedure of the metric formulation is described as follows.
@@
Given the above attention scores, we distinguish salient regions and define them as the Attention Focus. A threshold factor $\theta\in(0,1)$ is used to filter the significant values, resulting in a binarized Attention Focus Mask as follows:
$M_{A}=M(A_{w})=\textbf{1}\left\{A_{w}>\theta\cdot\max\left(A_{w}\right)\right\}\in\mathbb{R}^{N_{q}\times N_{k}},$ (6)
where $\textbf{1}\left\{\cdot\right\}$ represents an indicator function that returns 1 if the condition is true, otherwise 0.
@@
where the variables with a hat superscript ($\hat{A}_{w}$$\hat{M}_{A}$) denote the scores/masks from the quantized model. This metric lies in the range (0,1), and the distance function for PCC is defined as:
@@
Then, we leverage the defined Attention Focus Overlap metric to determine the optimal clipping boundaries (i.e. $x_{low}$ and $x_{up}$) for QK activations. Unlike conventional methods limited by distributional matching, our proposed clipping metric operates at a semantic level, enabling more functionally effective clipping decisions. As shown in Figure 3, while MSE-based calibration causes severe attention degradation, our PCC achieves closer results to the FP model, efficiently preserves the semantic capabilities of the quantized attention module.